DIRC Operations Weekly Report

Period: Wednesday 2007/1/24 - Wednesday 2007/2/7

Oncall: Bradley Wogsland
Fast Monitoring: Bradley Wogsland, Jose Benitez

File location: $BFROOT/www/Detector/DIRC/WeeklyReports/report_20070207.html
Plot directory: $BFROOT/www/Detector/DIRC/WeeklyReports/Plots/report_20070207 (empty)

Contents


BaBar-PEP news

PEP peak lumi: 5.8 x 1033 (previous week: ~0)

BaBar integrated lumi: 2.3 fb-1 (previous week: ~0)

Run 6 PEP delivery: 2.4 fb-1 (previous week: ~0)

Run 6 total delivery: 2.4 fb-1 Run 6 total BaBar recorded: 2.3 fb-1 including 0 fb-1 off-peak data

PEP activities this week:

  • ROD Jan 31. Chemical spill at MCC Feb 6.

    Foreseen work opportunities (PEP down periods):

  • no firm dates yet.

    BaBar activities/issues this week:

  • Vesda trip Feb 5. ROD Jan 31.

    DIRC activities

    See DIRC logbook for details

    Backgrounds

    Scheduled activities

    Pending Work

    Manpower situation

    Bradley and Jose rotating backup & on-call commissioning duties; Alejandro working doubletime soon.

    Links